If f (x) = (x^-x - x^0 + 2^x), find the following. (Enter your answer as a fraction or mixed number.)
f (3) =

Answer by Alan3354(69443)

If f (x) = (x^-x - x^0 + 2^x), find the following. (Enter your answer as a fraction or mixed number.)
f (3)
--------------
Sub 3 for each x
f(x) = x^-x - x^0 + 2^x
f(3) = 3^-3 - 3^0 + 2^3
= 1/27 - 1 + 8
= 7 + 1/27
= 190/27